# How to Use Large Language Models for Text Coding: The Case of Fatherhood Roles in Public Policy Documents

### v2

@@ 5.1 Quality
Interestingly, GPT-3 performs similarly when labeling the three tasks jointly and separately, in different runs. Performance on the second task is slightly better with joint labeling, while the contrary is true on the third task. This is good news because jointly labeling across the three tasks is faster and cheaper: we only need to prompt the model to label each text once instead of three times.
GPT-4
@@
When prompting GPT-4 with the best prompting strategy found in our experiments, it consistently outperforms GPT-3 and all individual human annotators. The fact that GPT-4 agrees more with individual human annotators than they do among themselves might appear counter-intuitive. However, we should remember that the models are coding the texts based on 15 examples whose labels all the human annotators agreed upon. Therefore, the models’ coding is heavily influenced by coding choices that are shared by annotators and reflect this agreement.
5.2 Speed
To evaluate the speed and cost of automatic coding, we let GPT-3777We choose GPT-3 over GPT-4 for this evaluation because it has higher rate limits. At the time of writing, GPT-4 was a brand-new technology, and its production was not yet mature. Therefore, GPT-3 was more representative of the speed and costs that the researchers will incur in the near future, when GPT-4 is expected to expand rate limits and lower costs. annotate our entire dataset of 1,910 sentences with the best prompting strategy resulting from the previous evaluation step, i.e., using 15-shot learning, providing a long, comprehensive description of the labels, and prompting the model to annotate across the three tasks jointly.
5.3 Cost
6 Discussion
Our results show that using an LLM for text coding is considerably faster and cheaper than recruiting, training, and evaluating human coders. It also achieves matching predictive performance and, in several cases, even performs better than individual human coders. Moreover, like human annotators, the LLM can annotate texts along multiple coding tasks without losing quality, and even improving it. We expect the outcome to be even better for English text coding and with languages that are more abundant online (such as Spanish or German), as they are more present in the language model training data openai2023gpt4 (32).
Of course, we can retrain human coders to improve their agreement with each other and their label accuracy, possibly outperforming LLM coding. However, that process requires even more time and incurs more costs.
But do we want to replace human coders with LLMs completely? That is a different question. LLMs are known to hallucinate (i.e., fabricate facts) and suffer from alignment issues (i.e., their moral underpinnings). Indeed, recent research has shown that different LLMs show different political leanings feng-etal-2023-pretraining (15). Hallucinations can be addressed by mapping non-matching labels to a default (for example, “not applicable” or the majority label), and repeated runs can reduce their risk. Moreover, the concern about hallucinations is larger when generating blocks of text. As we only request individual labels and give very explicit instructions, this problem is unlikely. Moral alignment is a trickier problem, although it can be solved for some tasks with adequate prompting argyle_out_2023 (5). These facts together make LLMs a sensible choice for sensitive data and for morally ambiguous or politically charged tasks. In these cases, constructing a validation set is suggested even more strongly to verify the alignment of LLM answers with human coding before deciding whether to annotate more data.
Labeling by LLM also offers another possibility, namely freeing up (human) time and resources to check the reliability of the coding. Too often, we assume that trained human coders are reliable and consistent, despite a growing body of evidence that this is not the case abercrombie-etal-2023-temporal (1). However, by outsourcing the first round of labeling to an LLM, we can use the time otherwise devoted to coding to check for inconsistencies, interrogate differences, and perform additional analyses.
@@ 7 Conclusion and Outlook
“If you want something done right, do it yourself.” For the longest time, this statement has highlighted the dilemma between efficient use of resources on the one hand and quality on the other. For a while now, the prospect of AI research assistants has promised to resolve this dilemma. But despite high quality, the technical expertise and vast training data requirements to operate these models have constituted severe thresholds for delegating to computers—delivering on promises of quality but eroding the prospects of resource efficiency. But as LLMs have been deployed, exhibiting outstanding results in cognitive tasks from economic research, writing, and data analysis (korinek_language_2023, 29, ) to acing the U.S. Bar exam and the Law School Admission Test (LSAT), they begged the question: If we want something done not only right but quickly and cheaply, is it finally time we let the computers do it?
In this paper, we apply LLMs to a typical downstream research task: data annotation. We present a case study on using LLMs (here, GPT-3 and GPT-4) to automate text coding, provide a general framework that can be applied to any text classification task, and discuss essential parameters and their effects on the outcome. Finally, we evaluate the quality of human-generated annotation compared to computer-generated annotation.
Using LLMs for text coding has several advantages over the traditional approach via either trained coders or supervised classifiers. Using LLMs is considerably faster (in our experiments, more than 270 times) and cheaper (at least 4 times) than relying on human coders. Regarding quality, GPT-3 (the cheaper alternative) performs on par with the individual human coders, while GPT-4 outperforms them. Though high quality and similar time gains can be achieved using supervised BERT-based classification models laurer_van2023_trans (31), LLMs require no specialized programming, can be implemented relatively simply via an API, do not require thousands of pre-coded examples to train the model, and can work with multiple different languages. By matching the results of individual human coders on a Swedish dataset with a seed set of just 15 coded examples, we show that LLMs are valuable for political scientists interested in text analysis in many languages.
An essential issue when working with LLMs is their prompting. This paper investigates the thoroughness of label definitions, the number of shots and the number of tasks included in the prompt. We find that all three aspects affect the performance of the LLMs. Few-shot learning has the most significant effect, but label definitions also improve performance. We also show that while annotating several tasks simultaneously has no clear reliability improvements, neither is the reliability lowered. This finding suggests that multiple interdependent annotation assignments can be performed at the same time and price by the LLMs without a reduction in quality but with an increase in cost-efficiency.
While our case study has limited scope, it adds to the growing body of evidence around LLMs as useful annotators. Our results suggest that if we want something done right, we might as well have an LLM do it.
@@
LLMs will likely have an outsized impact on political science methodology, not just in text analysis but also in text generation, such as survey questions, use of agent-based design, and manuscript outlines. The technology is here to stay, and we hope this paper helps practitioners to make sense of LLMs and adopt them responsibly and effectively for their needs.
@@ A.1 Probabilistic Language Models
Probabilistic language models of this type can be implemented in a few lines of code and assign reasonable probability estimates. Increasing the size of the $n$-grams produces better estimates. Still, it requires larger and larger corpora and more sophisticated smoothing chen1999empirical (9).
@@ A.2 Neural Language Models
Probabilistic language models were limited by their reliance on $n$-grams and the Markov horizon to compute the necessary statistics. Larger $n$-gram windows produced better models but required larger corpora and more storage and improved only incrementally. In the 2010s, a resurgence in the use of neural networks allowed for a dramatic leap forward in language models. Neural networks had been around since the 1950s rosenblatt1958perceptron (39), but interest in them died off in mainstream AI as they were too inconsistent in their results. A variety of factors led to their resurgence.
@@
Concurrently, researchers have made progress in understanding how these parameters should be computed and controlled, reducing the variability in neural network results. The development of better initialization regimes glorot2010understanding (16), better non-linear activation functions glorot2011deep (17), and an improved understanding of exploding or vanishing gradients hochreiter1991untersuchungen (24) has made it possible to train neural networks more efficiently and reliably.
The final factor that led to the resurgence of neural networks was computing power. Moore’s law schaller1997moore (41) predicted doubling the number of transistors on a microchip every two years. While that relation has slowed, the computing power available on a simple smartphone today is magnitudes larger than anything available at the dawn of AI and even as recently as 20 years ago. Computations that would have taken years to finish on a specialized supercomputer can now be run on a standard laptop. Specialized hardware like graphical processing units (GPUs) has become even more specialized and is the architectural backbone of today’s neural network models.
@@ A.3 The Transformer Architecture and GPT
Until recently, neural networks struggled with long-range dependencies, such as discourse phenomena in text sequences. In the sentence “This is not, by any stretch of the imagination, a funny movie,” the keywords not and funny are far apart, yet need to be connected to make sense of the whole. In 2017, (vaswani2017attention, 45, ) introduced a groundbreaking neural architecture called Transformer that surpassed previous architectures used for sequence modeling, such as recurrent hochreiter1997long (25, 10) and convolutional kim-2014-convolutional (28, 11) neural networks. At the heart of the Transformer architecture is the attention mechanism bahdanau2014neural (6), a neural module allowing the elements of a sequence to be linked directly to one another, even when distant. Transformer-based language models are today state-of-the-art and have gained public attention thanks to the remarkable linguistic prowess displayed by GPT models.
@@ Appendix C Prompts
Long labels definition: exaclty the same reported in the previous appendix (B).
@@ Appendix D Order of example sentences
Not only does the number of examples that we provide in the prompt affect the output of a language model, it also affects the order of the examples. The last examples in our 15-shot sample have a more significant effect on the model predictions than earlier sentences. This known the “recency effect” zhao2021calibrate (47) and is based on how the language model was trained. When predicting a missing word, the immediate context before the predicted word is more important than sentences that occurred several paragraphs before. In the following, we study the effect of this recency bias on labelling performance and a strategy to deal with it.

### v3

@@ 4.0.1 Prompt Engineering
Interestingly, GPT-3 performs similarly when labeling the three tasks jointly (multi-task) and separately, in different runs. Performance on the second task is slightly better with joint labeling, while the contrary is true on the third task. This is good news because we only need to prompt the model to label each text once instead of three times - reducing the time and monetary costs to almost a third.
4.1 Examples Order
Examples order Kappa Raw F1
0 49.36 62.57 54.92
1 48.27 61.90 53.51
2 46.58 60.67 53.72
Table 1: GPT-3’s average agreement with human coders on task 1, for three different random orders of the 15 labelled sentences provided in the prompt as examples.
Not only does the number of examples we provide in the prompt affect the output of a language model, but also the order of the examples, because LLMs are affected by a ‘recency bias” (zhao2021calibrate). Here, we study its effect on labelling performance. Table 2 shows the results of three different GPT-3 runs over the validation set, where the prompt contains a different random order of the 15 examples provided in each run. We only evaluate the most challenging task (Task 1: “type of paternal involvement”). We prompted the model with only the label list for this task and long, comprehensive descriptions. As expected from previous findings on recency bias, the performance changes with the order of the examples. Since we know which order performs best from trying different orders on the validation set, we can use this order within all other experiments involving few-shot learning.
4.2 Other Models
We then evaluate a mix of the most powerful models available to date using the best prompting strategy identified above (providing the full codebook).
@@ 4.2 Other Models
Noticeably, GPT-4 is on par with human annotators across the board. The fact that some LLMs agree with human annotators as much as humans agree between themselves (or even more) might appear counter-intuitive. However, we should remember that the models are coding the texts based on 15 examples whose labels all the human annotators agreed upon. Therefore, the models’ coding is heavily influenced by coding choices that are shared by annotators and reflect this agreement.
OS LLMs (GPT-SW3, Llama-2-13b and Mixtral-8x7b) underperform the proprietary (GPT) models on the first task. Interestingly, however, Llama-2-13b and Mixtral-8x7b catch up significantly on tasks 2 and 3, which are simpler (binary classification with less complex labels), becoming a viable alternative.
Our results show that the optimal coding choice for our tasks involves prompting the model with an exhaustive codebook, as the one provided to human annotators, and using GPT-4 in the multi-task setting or, in case of large datasets where GPT costs would be excessive, an OS LLM, which can be very competitive. For more details and evaluation metrics (raw agreement and Cohen’s $\kappa$) on our experiments, please refer to the Appendices. Clearly, the performance evaluated on our validation set may not be perfectly generalizable to the corpus as a whole. Ideally, subsequent manual evaluations could be conducted on a distinct set of sentences randomly selected from the rest of the corpus after its automatic annotation, performed with the optimal setup identified on the validation set.
@@ 5 Conclusion and Outlook
Many previous findings suggest that LLMs have great potential as annotators in the Political Sciences (ollion2023chatgpt). In this paper, we contribute to that strand of research by showing that, if prompted as human annotators, both proprietary and open-source LLMs can perform on par with humans even when the annotation involves non-English texts, legalese and political jargon, and complex labels based on abstract constructs. This complex setup, which is one encountered in many annotation projects for social science research, characterizes our first contribution with respect to the concurrent literature. Second, we show that, like human annotators, LLMs annotate best when provided with a detailed codebook that includes clear instructions, an exhaustive definition of the labels, and some labeled examples covering a variety of cases.
It may seem provocative that LLMs’ labels can have a higher quality than humans’. Most times, we treat human performance as the ground truth and the ability to mimic this performance as a measure of quality. As is the case in this paper. But as the far-from-perfect agreement between the humans make evident: humans make errors. While the errors may be small in relation to the closest coder, they might grow when comparing to the second closest, resulting in a subpar agreement on average. Of course, the LLMs make the same kind of errors, but it seems to distribute these errors so that it is closer to as many coders as possible, especially when all coders agree on the labelling of the few-shot examples provided to the model. In this way, the LLM becomes a more reliable mirror image of the average human annotation, than that of other humans.
Meanwhile, LLMs are much faster and cheaper than human coders. We estimated that, as of November 2023, GPT-3 could code our full corpus 270 times faster than what humans would, saving 60% of the money. [^4] For details on this estimate, see E.1. Our empirical results suggest that some OS LLMs are a viable alternative to proprietary models like GPT for very large corpora, which would incur intolerable costs even with proprietary models. In particular, we showed that OS LLMs can perform as well as the GPTs on less theoretically challenging tasks. In the near future, however, the performance of OS LLMs is likely to approach that of current proprietary models in every task.
@@
LLMs will likely have an outsized impact on political science methodology, not just in text analysis but also in text generation, such as survey questions, use of agent-based design, and manuscript outlines. The technology is here to stay, and we hope this paper helps practitioners to make sense of LLMs and adopt them responsibly and effectively for their needs. The software for evaluating and performing the annotation tasks of this paper was developed for seamless integration into other text coding projects. The script is publicly available at https://github.com/lorelupo/pappa.
@@ A.1 Probabilistic Language Models
Probabilistic language models of this type can be implemented in a few lines of code and assign reasonable probability estimates. Increasing the size of the $n$-grams produces better estimates. Still, it requires larger and larger corpora and more sophisticated smoothing (chen1999empirical).
@@ A.2 Neural Language Models
Probabilistic language models were limited by their reliance on $n$-grams and the Markov horizon to compute the necessary statistics. Larger $n$-gram windows produced better models but required larger corpora and more storage and improved only incrementally. In the 2010s, a resurgence in the use of neural networks allowed for a dramatic leap forward in language models. Neural networks had been around since the 1950s (rosenblatt1958perceptron), but interest in them died off in mainstream AI as they were too inconsistent in their results. A variety of factors led to their resurgence.
@@
Concurrently, researchers have made progress in understanding how these parameters should be computed and controlled, reducing the variability in neural network results. The development of better initialization regimes (glorot2010understanding), better non-linear activation functions (glorot2011deep), and an improved understanding of exploding or vanishing gradients (hochreiter1991untersuchungen) has made it possible to train neural networks more efficiently and reliably.
The final factor that led to the resurgence of neural networks was computing power. Moore’s law (schaller1997moore) predicted doubling the number of transistors on a microchip every two years. While that relation has slowed, the computing power available on a simple smartphone today is magnitudes larger than anything available at the dawn of AI and even as recently as 20 years ago. Computations that would have taken years to finish on a specialized supercomputer can now be run on a standard laptop. Specialized hardware like graphical processing units (GPUs) has become even more specialized and is the architectural backbone of today’s neural network models.
@@ A.3 The Transformer Architecture and GPT
Until recently, neural networks struggled with long-range dependencies, such as discourse phenomena in text sequences. In the sentence “This is not, by any stretch of the imagination, a funny movie,” the keywords not and funny are far apart, yet need to be connected to make sense of the whole. In 2017, vaswani2017attention introduced a groundbreaking neural architecture called Transformer that surpassed previous architectures used for sequence modeling, such as recurrent (hochreiter1997long; chung2014empirical) and convolutional (kim-2014-convolutional; dauphin2017language) neural networks. At the heart of the Transformer architecture is the attention mechanism (bahdanau2014neural), a neural module allowing the elements of a sequence to be linked directly to one another, even when distant. Transformer-based language models are today state-of-the-art and have gained public attention thanks to the remarkable linguistic prowess displayed by GPT models.
@@ Appendix C Prompts
Long labels definition: exactly the same reported in the previous appendix (B).
@@ Appendix D Order of example sentences
Not only does the number of examples that we provide in the prompt affect the output of a language model, also the order of the examples affect the output. The last examples in our 15-shot sample have a more significant effect on the model predictions than earlier sentences. This known the “recency effect” (zhao2021calibrate) and is based on how the language model was trained. When predicting a missing word, the immediate context before the predicted word is more important than sentences that occurred several paragraphs before. In the following, we study the effect of this recency bias on labelling performance and a strategy to deal with it.
@@
Appendix E Evaluation
To evaluate the speed and cost of automatic coding, we let GPT-3777We choose GPT-3 over GPT-4 for this evaluation because it has higher rate limits. At the time of writing, GPT-4 was a brand-new technology, and its production was not yet mature. Therefore, GPT-3 was more representative of the speed and costs that the researchers will incur in the near future, when GPT-4 is expected to expand rate limits and lower costs. annotate our entire dataset of 1,910 sentences with the best prompting strategy resulting from the previous evaluation step: 15-shot learning, providing a long, comprehensive description of the labels, and prompting the model to annotate across the three tasks jointly.
E.1 Speed
E.2 Cost
E.3 Quality
Down below follows more evaluation metrics on the dataset. Besides average F1-score, the average Cohens kappa and average raw agreement across each evaluation run is presented.
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 49.30 61.29 57.22
Human 2 n.a. n.a. n.a. 51.02 63.86 57.60
Human 3 n.a. n.a. n.a. 51.65 64.29 57.54
Humans AVG n.a. n.a. n.a. 50.66 63.15 57.45
GPT-3 1 none 0 13.41 27.71 22.73
GPT-3 1 short 0 20.89 36.38 34.27
GPT-3 1 long 0 18.06 33.90 34.69
GPT-3 1 none 15 36.53 51.05 44.22
GPT-3 1 short 15 44.50 58.48 49.40
GPT-3 1 long 15 49.36 62.57 54.92
GPT-3 1 reasoned 15 49.50 62.48 55.18
GPT-3 1,2,3 long 15 48.34 62.48 54.26
GPT-4 1,2,3 long 15 54.05 66.38 61.70
GPT-3 1 long 15 48.40 61.90 54.31
GPT-3 1 reasoned 15 48.93 62.19 54.49
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 26.74 54.86 50.13
Human 2 n.a. n.a. n.a. 26.55 58.29 48.61
Human 3 n.a. n.a. n.a. 29.75 59.14 52.33
Humans AVG n.a. n.a. n.a. 27.68 57.43 50.36
GPT-3 2 none 0 14.66 42.76 34.72
GPT-3 2 none 15 22.91 53.14 48.54
GPT-3 1,2,3 none 15 25.91 59.14 50.07
GPT-4 1,2,3 none 15 31.97 63.62 53.49
Annotator Tasks Label description N. Examples Kappa Raw F1
Human 1 n.a. n.a. n.a. 41.64 66.57 61.83
Human 2 n.a. n.a. n.a. 37.00 67.14 57.94
Human 3 n.a. n.a. n.a. 44.17 71.14 63.20
Humans AVG n.a. n.a. n.a. 40.94 68.28 60.99
GPT-3 3 none 0 7.35 40.29 29.94
GPT-3 3 short 0 7.96 36.86 32.00
GPT-3 3 none 15 39.75 68.76 60.97
GPT-3 3 short 15 44.08 70.86 64.25
GPT-3 1,2,3 short 15 41.85 70.48 62.67
GPT-4 1,2,3 short 15 40.55 71.81 61.43